# DEEPTalk: Dynamic Emotion Embedding for Probabilistic Speech-Driven 3D Face Animation

### v1

@@ Abstract
Speech-driven 3D facial animation has garnered lots of attention thanks to its broad range of applications. Despite recent advancements in achieving realistic lip motion, current methods fail to capture the nuanced emotional undertones conveyed through speech and produce monotonous facial motion. These limitations result in blunt and repetitive facial animations, reducing user engagement and hindering their applicability. To address these challenges, we introduce DEEPTalk, a novel approach that generates diverse and emotionally rich 3D facial expressions directly from speech inputs. To achieve this, we first train DEE (Dynamic Emotion Embedding), which employs probabilistic contrastive learning to forge a joint emotion embedding space for both speech and facial motion. This probabilistic framework captures the uncertainty in interpreting emotions from speech and facial motion, enabling the derivation of emotion vectors from its multifaceted space. Moreover, to generate dynamic facial motion, we design TH-VQVAE (Temporally Hierarchical VQ-VAE) as an expressive and robust motion prior overcoming limitations of VAEs and VQ-VAEs. Utilizing these strong priors, we develop DEEPTalk, A talking head generator that non-autoregressively predicts codebook indices to create dynamic facial motion, incorporating a novel emotion consistency loss. Extensive experiments on various datasets demonstrate the effectiveness of our approach in creating diverse, emotionally expressive talking faces that maintain accurate lip-sync. Source code will be made publicly available soon.
footnotetext: * These authors contributed equally to this work.footnotetext: Corresponding author
@@ Introduction
Previous studies on talking faces have generated expressions using either emotion labels (Daněček et al. 2023; Gan et al. 2023; Pan et al. 2023; Ji et al. 2021) or reference expressions (Ji et al. 2022; Tan et al. 2024; Ma et al. 2023a; Liang et al. 2022). Using emotion labels provides expressive but limited outcomes, while reference expressions offer more diversity at the cost of needing numerous expressive references. Moreover, both approaches fail to capture vocal nuances, often leading to misalignment between facial expressions and speech. As shown in Figure 1, the phrase ”That’s just what I needed!” can be expressed with different emotions (e.g., happiness or anger) based on its delivery, highlighting the need for facial expressions that match the intended meaning. Without this alignment, expressions may appear unnaturally disconnected from the spoken words, contributing to the uncanny valley effect (Wang, Lilienfeld, and Rochat 2015).
@@
We then aim to build an expressive motion prior that is robust to perceptual losses. Recent studies have demonstrated that by learning motion priors from discrete codebooks, it is possible to generate a diverse range of facial and body motions (Li et al. 2021; Yi et al. 2023; Xing et al. 2023; Ng et al. 2022). However, due to the dynamic nature of emotional talking faces, VQ-VAE(Van Den Oord, Vinyals et al. 2017) alone struggles to capture the entire motion space fully. Recognizing that facial motion encompasses different temporal hierarchies—for instance, the mouth region exhibits high frequencies while the upper face displays lower temporal frequencies—we train a hierarchical discrete motion prior to addressing these variations effectively.
@@
  • We propose a novel temporally hierarchical VQVAE (TH-VQVAE) to build a motion prior that is both expressive and robust.
@@
The availability of large datasets (Richard et al. 2021; Fanelli et al. 2010), comprising 4D meshes synchronized with audio, has significantly advanced the development of deep learning-based methods, which exhibit robust lip synchronization on 3D faces. Extending these advancements, VOCA (Cudeiro et al. 2019) improves realism by generating animations from any speech inputs, employing time convolutions along with a speaker identity one-hot vector. FaceFormer(Fan et al. 2022) employs a transformer-based model to generate facial movements auto-regressively. However, despite their accurate lip synchronization, the generated upper face remains static as they are less correlated with the input speech. MeshTalk (Richard et al. 2021) addresses this issue by disentangling audio-correlated and uncorrelated movements using a categorical latent space to model upper face dynamics. Similarly, CodeTalker (Xing et al. 2023) utilizes a discrete motion prior with VQ-VAE to reconstruct dynamic facial motions across the entire face. However, these methods generate deterministic motion, overlooking the inherently non-deterministic relationship between facial motion and speech. Therefore, FaceDiffuser (Stan, Haque, and Yumak 2023) addresses the limitation of deterministic models by employing a diffusion model, which is probabilistic in nature, to generate multiple facial motions for a given speech. However, it still falls short in capturing the diverse emotional expressions conveyed through speech, which is crucial for enhancing interactiveness for talking face models.
@@
Recent studies have underscored the critical role of emotion in creating realistic and expressive 3D facial motions by integrating emotional information to enhance facial animation. Specifically, EMOTE (Daněček et al. 2023) employs one-hot labels to control emotion, producing emotional facial motions through an emotion-content disentanglement loss. Chen et al. (Chen, Zhao, and Zhang 2023) utilized the logits from an emotion classifier applied to reference images as an emotion prior during training, generating emotional facial motion through an emotion-augmented network. However, these methods require explicit control, lacking a direct connection to the emotion conveyed in the actual speech. The work most related to ours is EmoTalk (Peng et al. 2023), which aims to generate emotional blend shapes from audio input only, utilizing an emotion-content disentangling method. However, the training of EmoTalk’s emotion-content disentangling encoder is constrained to specific datasets like RAVDESS(Livingstone and Russo 2018), containing the same sentences expressed in multiple emotions. This constraint limits the use of larger, more diverse datasets such as MEAD(Wang et al. 2020), ultimately hindering the model’s ability to generalize emotional expressions on other datasets. Unlike previous methods that either disentangle content and emotion from speech or rely on explicit conditioning, our approach establishes a robust emotion prior by leveraging the natural correlation between speech and facial expressions, enabling seamless integration into the talking face training pipeline.
@@
Our goal is to synthesize emotional facial motion solely from speech. However, the complex many-to-many relationship between speech and facial expressions poses challenges for generating expressive emotional faces. To overcome this, we build (1) DEE, a joint probabilistic emotional space to encode both audio and facial motion. DEE sample emotion embeddings from audio, and enable generated facial motions aligned to the corresponding embedding. Also for robust and expressive motion prior, we design (2) TH-VQVAE to learn a discrete motion space capable of capturing both high and low-frequency motions. Building on these priors, we train (3) DEEPTalk to map emotional audio to motion codebooks, resulting in facial animations that are both emotionally expressive and accurately lip-synced. Additional details on each component are provided in the supplementary material.
@@
Our goal is to analyze $\boldsymbol{A}_{1:T}$ on emotions using DEE encoder and contents using audio encoder along with one-hot speaker identity $\boldsymbol{c}_{id}$ to predict FLAME expression parameters $\boldsymbol{\hat{F}}_{1:T}=(\boldsymbol{\hat{f}}_{1},\cdots,\boldsymbol{\hat{f}}_{T})$. As our talking head model is non-auto-regressive, denoting $\theta$ as model parameters, our end-to-end procedure can be written as
@@
We utilized the recently proposed emotion2vec (Ma et al. 2023b) to extract emotion features in our audio encoder. Denoting feature extractor as $F$, this process is defined as :
$F^{emo}_{audio}(\boldsymbol{A}_{1:T})\rightarrow\epsilon_{audio}$ (3)
For the expression feature extractor in the expression encoder, we first trained an emotion classification model using the AffectNet dataset (Mollahosseini, Hasani, and Mahoor 2017). As Affectnet is an image dataset, we applied a 3D flame parameter reconstruction method (Daněček, Black, and Bolkart 2022) to generate 3D pseudo ground truth for each image in AffectNet and trained an emotion recognition model. The trained encoder was then repurposed as the facial expression feature extractor $F$. This process is defined as:
$F^{emo}_{exp}(\boldsymbol{F}_{1:t})\rightarrow\epsilon_{exp}$ (4)
@@
With both audio and expression feature extractors in place, DEE trains the audio and expression encoders separately. The overview of DEE is illustrated in Figure 2(a). Each encoder $E$ comprises two distinct heads, followed by a Generalized Pooling Operator (GPO) (Chen et al. 2021), with one head dedicated to $\mu$ and the other to $\log\sigma^{2}$, similar to the approach in PCME++(Chun 2023). Therefore, the final probabilistic emotion embeddings for audio and expression are as follows :
@@
FID $\downarrow$ FFD $\downarrow$ Emo-FID $\downarrow$ LSE-D $\downarrow$ LSE-C $\uparrow$ FID $\downarrow$ FFD $\downarrow$ Emo-FID $\downarrow$ LSE-D $\downarrow$ LSE-C $\uparrow$ FID $\downarrow$ FFD $\downarrow$ Emo-FID $\downarrow$ LSE-D $\downarrow$ LSE-C $\uparrow$
@@
We computed the Mean Squared Error between the pseudo-GT and predicted vertices.
@@
To ensure that the generated face motion reflects the same emotion as the input audio, we proposed an emotion consistency loss. By leveraging the trained DEE, we sampled DEE embedding from audio input $e_{a}$ and the generated motion $e_{exp}$ and enforced the high cosine similarity between these pairs to ensure they represent the same emotion.
$L_{emo}=\frac{{e_{a}\cdot e_{exp}}}{{\|e_{a}\|\|e_{exp}\|}}$ (16)
@@
We adopt FID to evaluate the realism of rendered faces. To further assess the realism of facial movements in the FLAME parameter space, we developed FFD (Frechet Face Distance), inspired by FGD (Yoon et al. 2020), and computed it on sequences of FLAME parameters using an encoder trained on MEAD and BEATv2 (Liu et al. 2024). To evaluate emotional expressiveness, we compute Emo-FID, an adaptation of FID that uses an emotion feature extractor from AffectNet (Mollahosseini, Hasani, and Mahoor 2017), replacing the inception network to focus on emotion properties. For lip-sync evaluation, we used the SyncNet score (Chung and Zisserman 2016) as Lip Vertex Error (LVE) is unsuitable for evaluating DEEPTalk’s lip generation performance. Since DEEPTalk generates diverse emotional faces, the lip vertices between the ground truth and predictions may differ even when syllables are correctly pronounced. Therefore, we employed LSE-D (Lip Sync Error Distance) and LSE-C (Lip Sync Error Confidence) from Syncnet as our evaluation metrics following (Aneja et al. 2024). Finally, we measure Diversity (Ren et al. 2023).
@@
In Table 1, DEEPTalk achieves the highest performance on LSE-D and ranks first or second on LSE-C across the MEAD test and other datasets, demonstrating strong generalization and precise lip synchronization. Also we conducted evaluations on Emo-Vox for lip generalization and fair comparison, using both vertex and paramter-based models. Table 2 shows our model excels in LSE-D and ranks second in LSE-C, demonstrating effective generalization to emotional, in-the-wild audio. This performance further indicates our bottom codebook’s capability to capture high-frequency details enhancing dynamic lip movements. In contrast, even though being trained on the ground truth scan dataset, the vertex-based models fail to achieve accurate lip synchronization compared to those using pseudo ground truth. Note that FaceFormer, trained on MEAD, produces nearly neutral expressions (see Figure 3), while ours are more expressive, benefiting FaceFormer’s lip-sync performance due to the emotion and lip-sync trade off, as demonstrated in ablation studies.
Method $\alpha$ $\tau$ Diversity ($\uparrow$) LSE-D ($\downarrow$) LSE-C ($\uparrow$)
@@
DEEPTalk provides control over the diversity of generated facial motions in two ways: 1) diverse emotional facial motions from the same speech input by adjusting the control factor $\alpha$, and 2) diverse facial motions from the same speech and emotion by modifying the temperature of the bottom ($\tau_{b}$) and top ($\tau_{t}$) codebook predictors. We evaluated each controllability factor by varying them independently while keeping the other factor deterministic. We continued this process until our method surpassed FaceDiffuser in terms of diversity, after which we assessed lip synchronization. As shown in Table 3, DEEPTalk demonstrates superior lip synchronization even with increased diversity than FaceDiffuser on both control factors, illustrating the model’s effectiveness. Notably, these two methods of control are orthogonal and can be used independently, potentially leading to even greater diversity.
@@
Figure 4: TSNE Visualization of DEE. Each embeddings are clustered according to emotional categories.
@@
In Figure 3, we compare our method with SOTA methods on the MEAD test set. While most methods generate natural lip movements, they sometimes misalign with the ground truth, such as opening or closing lips incorrectly. Additionally, except for EMOTE, other methods often produce incorrect or unexpressive expressions, failing to match the ground truth. This highlights the effect of DEE and $L_{emo}$ on generating emotional faces. While EMOTE generates emotional expressions using emotional labels, it often produces exaggerated faces, like unnaturally wide eyes or flattened eyebrows, which deviate from the ground truth. In contrast, DEEPTalk generates natural emotional expressions that closely resemble the ground truth faces without the need for GT emotion labels by leveraging the emotions inherent within the speech.
@@
Figure 4 illustrates the clustering of our emotion embedding by emotions, demonstrating that it effectively captures a meaningful emotion space. Furthermore, to evaluate its efficacy in generating emotional expressions, we conducted an experiment in which we randomly selected two audio samples with distinct emotions and interchanged their emotion embeddings from DEE’s audio encoder, before inputting them into the Face Feature Extractor. As shown in Figure 6, this swap effectively modifies the emotional expression to match the reference speech, while maintaining precise lip synchronization with the original speech. Further analysis of our emotion embedding is provided in the Supplementary.
Figure 6: Swapping emotion embedding alters the emotional expression of the speech to align with the reference, while maintaining precise lip synchronization
@@
Methods LSE-D $\downarrow$ LSE-C $\uparrow$ LSE-D $\downarrow$ LSE-C $\uparrow$ Emo-FID $\downarrow$ FID $\downarrow$ FFD $\downarrow$

### v2

@@ Abstract
Speech-driven 3D facial animation has garnered lots of attention thanks to its broad range of applications. Despite recent advancements in achieving realistic lip motion, current methods fail to capture the nuanced emotional undertones conveyed through speech and produce monotonous facial motion. These limitations result in blunt and repetitive facial animations, reducing user engagement and hindering their applicability. To address these challenges, we introduce DEEPTalk, a novel approach that generates diverse and emotionally rich 3D facial expressions directly from speech inputs. To achieve this, we first train DEE (Dynamic Emotion Embedding), which employs probabilistic contrastive learning to forge a joint emotion embedding space for both speech and facial motion. This probabilistic framework captures the uncertainty in interpreting emotions from speech and facial motion, enabling the derivation of emotion vectors from its multifaceted space. Moreover, to generate dynamic facial motion, we design TH-VQVAE (Temporally Hierarchical VQ-VAE) as an expressive and robust motion prior overcoming limitations of VAEs and VQ-VAEs. Utilizing these strong priors, we develop DEEPTalk, a talking head generator that non-autoregressively predicts codebook indices to create dynamic facial motion, incorporating a novel emotion consistency loss. Extensive experiments on various datasets demonstrate the effectiveness of our approach in creating diverse, emotionally expressive talking faces that maintain accurate lip-sync. Our project page is available at: https://whwjdqls.github.io/deeptalk.github.io/.
@@ Introduction
Previous studies on talking faces have generated expressions using either emotion labels (Daněček et al. 2023; Gan et al. 2023; Pan et al. 2023; Ji et al. 2021) or reference expressions (Ji et al. 2022; Tan et al. 2024; Ma et al. 2023a; Liang et al. 2022). Using emotion labels provides expressive but limited outcomes, while reference expressions offer more diversity at the cost of needing numerous expressive references. Moreover, both approaches fail to capture vocal nuances, often leading to misalignment between facial expressions and speech. As shown in Figure 1, the phrase ”That’s just what I needed!” can carry different meanings based on the emotion it conveys (e.g., happiness or anger), emphasizing the importance of facial expressions that reflect the intended sentiment. Without this alignment, expressions may appear unnaturally disconnected from the spoken words, contributing to the uncanny valley effect (Wang, Lilienfeld, and Rochat 2015).
@@
We then aim to build an expressive motion prior that is robust to perceptual losses. Recent studies have demonstrated that by learning motion priors from discrete codebooks, it is possible to generate a diverse range of facial and body motions (Li et al. 2021; Yi et al. 2023; Xing et al. 2023; Ng et al. 2022). However, due to the dynamic nature of emotional talking faces, VQ-VAE(Van Den Oord, Vinyals et al. 2017) alone struggles to capture the entire motion space fully. Recognizing that facial motion encompasses different temporal hierarchies—for instance, the mouth region exhibits high frequencies while the upper face displays lower temporal frequencies—we train a hierarchical discrete motion prior to effectively address these variations effectively.
@@
  • We propose a novel temporally hierarchical VQ-VAE (TH-VQVAE) to build a motion prior that is both expressive and robust.
@@
The availability of large 4D mesh datasets synchronized with audio (Richard et al. 2021; Fanelli et al. 2010) has greatly advanced deep learning-based 3D face animation, enabling robust lip synchronization. Extending these advancements, VOCA (Cudeiro et al. 2019) improves realism by generating animations from any speech inputs, employing time convolutions along with a speaker identity one-hot vector. FaceFormer (Fan et al. 2022) employs a transformer-based model to generate facial movements auto-regressively. However, despite their accurate lip synchronization, the generated upper face remains static as they are less correlated with the input speech. MeshTalk (Richard et al. 2021) addresses this issue by disentangling audio-correlated and uncorrelated movements using a categorical latent space to model upper face dynamics. Similarly, CodeTalker (Xing et al. 2023) utilizes a discrete motion prior with VQ-VAE to reconstruct dynamic facial motions across the entire face. However, these methods generate deterministic motion, overlooking the inherently non-deterministic relationship between facial motion and speech. Therefore, FaceDiffuser (Stan, Haque, and Yumak 2023) addresses the limitation of deterministic models by employing a diffusion model, which is probabilistic in nature, to generate multiple facial motions for a given speech. However, it still falls short in capturing the diverse emotional expressions conveyed through speech, which is crucial for enhancing interactiveness for talking face models.
@@
Recent studies have underscored the critical role of emotion in creating realistic and expressive 3D facial motions by integrating additional emotional information. Specifically, EMOTE (Daněček et al. 2023) employs one-hot labels to control emotion, producing emotional facial motions through an emotion-content disentanglement loss. Chen et al. (Chen, Zhao, and Zhang 2023) utilized the logits from an emotion classifier applied to reference images as an emotion prior during training, generating emotional facial motion through an emotion-augmented network. However, these methods require explicit control, lacking a direct connection to the emotion conveyed in the actual speech. The work most related to ours is EmoTalk (Peng et al. 2023), which aims to generate emotional blendshapes from audio input only, utilizing an emotion-content disentangling method. However, the training of EmoTalk’s emotion-content disentangling encoder is constrained to specific datasets like RAVDESS (Livingstone and Russo 2018), containing the same sentences expressed in multiple emotions. This constraint limits the use of larger, more diverse datasets such as MEAD (Wang et al. 2020), ultimately hindering the model’s ability to generalize emotional expressions on other datasets. Unlike previous methods that either disentangle content and emotion from speech or rely on explicit conditioning, our approach establishes a robust emotion prior by leveraging the natural correlation between speech and facial expressions, enabling seamless integration into the talking face training pipeline.
@@
Our goal is to synthesize emotional facial motion solely from speech. However, the complex many-to-many relationship between speech and facial expressions poses challenges for generating expressive emotional faces. To overcome this, we build (1) DEE, a joint probabilistic emotional space to encode both audio and facial motion. DEE sample emotion embeddings from audio, and enable generated facial motions aligned to the corresponding embedding. Also for robust and expressive motion prior, we design (2) TH-VQVAE to learn a discrete motion space capable of capturing both high and low-frequency motions. Building on these priors, we train (3) DEEPTalk to map emotional audio to motion codebooks, resulting in facial animations that are both emotionally expressive and accurately lip-synced. Additional details on each component are provided in the Supplementary.
@@
Our goal is to analyze the content and emotion of $\boldsymbol{A}_{1:T}$ and, together with the one-hot speaker identity $\boldsymbol{c}_{id}$, predict FLAME expression parameters $\boldsymbol{\hat{F}}_{1:T}=(\boldsymbol{\hat{f}}_{1},\cdots,\boldsymbol{\hat{f}}_{T})$ that align with the input speech. As our talking head model is non-auto-regressive, denoting $\theta$ as model parameters, our end-to-end procedure can be written as
@@
We utilized the recently proposed emotion2vec (Ma et al. 2023b) to extract emotion features in our audio encoder. Denoting feature extractor as $F_{audio}$, this process is defined as :
$F_{audio}(\boldsymbol{A}_{1:T})\rightarrow\epsilon_{audio}$ (3)
For the expression feature extractor in the expression encoder, we first trained an emotion classification model using the AffectNet dataset (Mollahosseini, Hasani, and Mahoor 2017). As Affectnet is an image dataset, we applied a 3D flame parameter reconstruction method (Daněček, Black, and Bolkart 2022) to generate 3D pseudo ground truth for each image in AffectNet and trained an emotion recognition model. The trained encoder was then repurposed as the facial expression feature extractor $F_{exp}$. This process is defined as:
$F_{exp}(\boldsymbol{F}_{1:T})\rightarrow\epsilon_{exp}$ (4)
@@
With both audio and expression feature extractors in place, DEE trains the audio and expression encoders separately. The overview of DEE is illustrated in Figure 2(a). Each encoder $E$ comprises two distinct heads, followed by a Generalized Pooling Operator (GPO) (Chen et al. 2021), with one head dedicated to $\mu$ and the other to $\log\sigma^{2}$, similar to the approach in PCME++(Chun 2023). The final probabilistic emotion embeddings for audio and expression can be written as follows:
@@
[…]
@@
We computed the Mean Squared Error between the pseudo-GT and predicted vertices, denoted as $L_{rec}$.
@@
To ensure that the generated face motion reflects the same emotion as the input audio, we propose an emotion consistency loss. By leveraging the trained DEE, we predict the mean from audio input $\mu_{a}$ and generated motion $\mu_{e}$ and enforce a high cosine similarity between these pairs to ensure they represent the same emotion.
$L_{emo}=\frac{{\mu_{a}\cdot\mu_{e}}}{{\|\mu_{a}\|\|\mu_{e}\|}}$ (16)
@@
We adopt FID to evaluate the realism of rendered faces. To further assess the realism of facial movements in the FLAME parameter space, we developed FFD (Frechet Face Distance), inspired by FGD (Yoon et al. 2020), and computed it on sequences of FLAME parameters using an encoder trained on MEAD and BEATv2 (Liu et al. 2024). To evaluate emotional expressiveness, we compute Emo-FID, an adaptation of FID that uses an emotion feature extractor from AffectNet (Mollahosseini, Hasani, and Mahoor 2017), replacing the inception network to focus on emotion properties. For lip-sync evaluation, we used the SyncNet score (Chung and Zisserman 2016) as Lip Vertex Error (LVE) is unsuitable for evaluating DEEPTalk’s lip generation performance. Since DEEPTalk generates diverse emotional faces, the lip vertices between the ground truth and predictions may differ even when syllables are correctly pronounced. Therefore, we employed LSE-D (Lip Sync Error Distance) and LSE-C (Lip Sync Error Confidence) from Syncnet as our evaluation metrics following (Aneja et al. 2024), as both metrics do not require ground truth lip movements. Finally, we measure Diversity (Ren et al. 2023).
@@
In Table 1, DEEPTalk achieves the highest performance on LSE-D and ranks first or second on LSE-C across the MEAD test set and other datasets, demonstrating strong generalization and precise lip synchronization. We also conducted evaluations on Emo-Vox to assess lip synchronization, while ensuring a fair comparison with both parameter-based and vertex-based models. As shown in Table 2, our model excels in LSE-D and ranks second in LSE-C, demonstrating effective generalization of lip movements to emotional, in-the-wild audio. This performance further indicates our bottom codebook’s capability to capture high-frequency details, enhancing dynamic lip movements. In contrast, despite being trained on the ground truth scan dataset, vertex-based models fail to achieve accurate lip synchronization compared to those using pseudo ground truth. Notably, FaceFormer, trained on MEAD, produces nearly neutral expressions (see Figure 3), whereas ours are more expressive, benefiting FaceFormer’s lip-sync performance due to the emotion and lip-sync trade off, as demonstrated in ablation studies.
[…]
@@
DEEPTalk provides control over the diversity of generated facial motions in two ways: 1) diverse emotional facial motions from the same speech input by adjusting the control factor $\alpha$, and 2) diverse facial motions from the same speech and emotion by modifying the temperature of the bottom ($\tau_{b}$) and top ($\tau_{t}$) codebook predictors. We evaluated each controllability factor by varying them independently while keeping the other factor deterministic. We continued this process until our method surpassed FaceDiffuser in terms of diversity, after which we assessed lip synchronization. As shown in Table 3, DEEPTalk demonstrates superior lip synchronization, even with increased diversity, compared to FaceDiffuser on both control factors, illustrating the model’s effectiveness in generating diverse yet accurate facial motion. Notably, these two methods of control are orthogonal and can be used independently, potentially leading to even greater diversity.
@@
Figure 4: Embeddings are clustered by emotion categories.
@@
In Figure 3, we compare our method with SOTA methods on the MEAD test set. While most methods generate natural lip movements, they sometimes misalign with the ground truth, such as opening or closing lips incorrectly. Additionally, except for EMOTE, other methods often produce incorrect or unexpressive expressions, failing to match the ground truth. This highlights the effect of DEE and $L_{emo}$ on our framework. While EMOTE generates emotional expressions using emotional labels, it often produces exaggerated faces, like unnaturally wide eyes or flattened eyebrows, which deviate from the ground truth. In contrast, DEEPTalk generates natural emotional expressions that closely resemble the ground truth faces without the need for GT emotion labels by leveraging the emotions inherent within the speech.
@@
Figure 4 illustrates the clustering of our emotion embedding by emotions, utilizing T-SNE, demonstrating that it effectively captures a meaningful emotion space. Furthermore, to evaluate its efficacy in generating emotional expressions, we conducted an experiment in which we randomly selected two audio samples with distinct emotions and interchanged their emotion embeddings from DEE’s audio encoder, before inputting them into the Face Feature Extractor. As shown in Figure 6, this swap effectively modifies the emotional expression to match the reference speech, while maintaining precise lip synchronization with the original speech. Further analysis of emotion embedding is provided in the Supplementary.
Figure 6: Swapping emotion embedding alters the emotional expression while maintaining precise lip synchronization
@@
[…]